A square metal plate with a thickness of 1.5 cm has no net charge and is placed in a region of uniform electric field 8.0 x 104 N/C directed perpendicularly to the plate. Find the resulting surface charge density on each face of the plate.


1
Expert's answer
2021-05-07T17:27:55-0400
"E=\frac{\sigma}{\epsilon_0}\\8\cdot10^{4}=\frac{\sigma}{8.85\cdot10^{-12}}\\
\sigma=7.1\cdot10^{-7}\frac{C}{m^2}"